OpenStudy (anonymous):

Solve lim h->0 h^3 + 3h^2 x + 3h x^2 + x^3 over h

OpenStudy (anonymous):

@RadEn can u help? :)

OpenStudy (anonymous):

differentiate the numerator and the denominator

OpenStudy (anonymous):

i expanded it wrong?

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ h^3+3x^2h+x^3 }{ h }\]

OpenStudy (anonymous):

can you confirm this is the equation you are asking for

OpenStudy (anonymous):

no its this one \[ \lim_{h \rightarrow 0} h^3 + 3h^2 x + 3h x^2 + x^3\]

OpenStudy (anonymous):

over h

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{ h^3 + 3h^2 x + 3h x^2 + x^3 }{ h }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because anything/0 is undefined in your case then i will apply the famous L'Hopital rule, which allows you to differentiate the numerator and the denominator.

OpenStudy (anonymous):

ok how do you do that?

OpenStudy (anonymous):

i will show you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0}\frac{ 3h^2+6xh^2+3x^2 }{ 1 }\]

OpenStudy (anonymous):

so when you plug the zero, you are left with 3x^2

OpenStudy (anonymous):

remember we differentiating with respect to h

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

x would be a constant in this case

OpenStudy (anonymous):

so only 2x^2 will remain

OpenStudy (anonymous):

3x^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

you are welcome! all the best

hartnn (hartnn):

L'Hopitals can be applied only if the form is 0/0 or infinity / infinity , here its not the case, so L'Hopital's can't be applied.

OpenStudy (anonymous):

ohh so how would i do it then?

hartnn (hartnn):

are you sure this: h^3 + 3h^2 x + 3h x^2 + x^3 is the numerator ? or was it (x+h)^3 -x^3 ??

OpenStudy (anonymous):

the original question was with Compute f’(a) using the limit definition f(x) = x^3 + 4, a = -1, so expanded (x + h)^3 and I got that one.

hartnn (hartnn):

aha! do you know the limit definition of derivative ?

OpenStudy (anonymous):

no

hartnn (hartnn):

let me tell you :) \(\large f'(x)=\lim \limits_{h \rightarrow 0} \dfrac{f(x+h)-f(x)}{h}\) you computed f(x+h) = (x+h)^3 +4 but missed the '-f(x)' part of numerator.

OpenStudy (anonymous):

ohhh you're right!

hartnn (hartnn):

so, now your numerator will become ?

OpenStudy (anonymous):

(x + h)^3 - 4?

hartnn (hartnn):

lets go term by term : f(x) = x^3+4 so what will be f(x+h) =...? do you know how to find it ?

OpenStudy (anonymous):

no

hartnn (hartnn):

to get f(x+h), just replace 'x' in f(x) by '(x+h)' so what u get by replacing 'x' in x^3+4 by 'x+h' ?

OpenStudy (anonymous):

oh so (x +h)^3 + 4?

hartnn (hartnn):

yes, thats f(x+h) now your numerator, f(x+h)-f(x) = (x+h)^3 +4 - (x^3+4) right ? can you simplify the numerator ?

OpenStudy (anonymous):

ok hold on

hartnn (hartnn):

sure, take your time :)

OpenStudy (anonymous):

so i got h^3 + 3h^2 x + 3h x^2 ?

hartnn (hartnn):

That.is.absolutely.correct! Good :) now comes the easy part, so, your limit now is ?

OpenStudy (anonymous):

okay hold on lol so i just cross the like terms right?

OpenStudy (anonymous):

so i got h^3 over h ?

hartnn (hartnn):

cross ? can you factor out h from numerator ?

OpenStudy (anonymous):

ohh okay

hartnn (hartnn):

wait, your limit is : \(\large f'(x)=\lim \limits_{h \rightarrow 0} \dfrac{h^3 + 3h^2 x + 3h x^2 }{h}\) right ? factor out 'h' from numerator.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if i do that only h3 will remain right?

hartnn (hartnn):

umm, no... factor out one 'h' from each term of numerator, like this : \(h^3+3h^2x+3hx^2= h (h^2+3hx+3x^2)\) got this ?

OpenStudy (anonymous):

ohh ok

hartnn (hartnn):

now you notice that this factored out 'h' from numerator, gets cancelled with 'h' in denominator ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, your limit now is : \(\large \lim \limits_{h \rightarrow 0}[h^2+3hx+3x^2]\) right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

now you can just plug in h=0 in that to get the value of limit, what u get after plgging in h=0 ?

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

ok so only h remains

hartnn (hartnn):

h ? you are putting h=0

OpenStudy (anonymous):

you meant put zero in every h right?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

yea thats what i did

OpenStudy (anonymous):

wait never mind

OpenStudy (anonymous):

nothind remains

OpenStudy (anonymous):

they all equal to zeros

hartnn (hartnn):

keep 'x' term as it is,

OpenStudy (anonymous):

oh ok

hartnn (hartnn):